# On the number of modes of Gaussian kernel density estimators

### v2

@@ Appendix A Additional proofs
$\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|\lesssim\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}\hskip 28.45274pt\text{ as }\xi_{1},\xi_{2}\to\infty.$
Observe that the critical points of $g$ are ±$\pm\sqrt{\frac{1}{\beta}}$, whereas those of $g^{\prime}$ are 00 and ±$\pm\sqrt{\frac{3}{\beta}}$. Since $\beta\to\infty$ as well, pick $\varepsilon>0$ sufficiently small and such that $\varepsilon>\sqrt{\frac{3}{\beta}}$. We first see that
$\mathopen{}\mathclose{{}\left|\int_{|x|>\varepsilon}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$=\mathopen{}\mathclose{{}\left|\int_{|x|>\varepsilon}\frac{1}{\mathrm{i}}\frac{1}{\xi_{1}g^{\prime}(x)+\xi_{2}g^{\prime\prime}(x)}\frac{\,\mathrm{d}}{\,\mathrm{d}x}\mathopen{}\mathclose{{}\left(e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}}\right)e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$\lesssim\frac{1}{\xi_{1}+\xi_{2}},$
where we used integration by parts to obtain the last inequality—this is in fact an elementary version of the method of non-stationary phase. For the integral over $\{|x|\leqslant\varepsilon\}$, we look to use the method of stationary phase as $\xi_{1},\xi_{2}\to\infty$, by distinguishing the three regimes $\xi_{1}\gg\xi_{2}$, $\xi_{2}\gg\xi_{1}$, and $\xi_{1}\sim\xi_{2}$. When $\xi_{1}\gg\xi_{2}$, we have
@@
$\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
@@
$=\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}+\xi_{2})\mathopen{}\mathclose{{}\left(\frac{\xi_{1}}{\xi_{1}+\xi_{2}}g(x)+\frac{\xi_{2}}{\xi_{1}+\xi_{2}}g^{\prime}(x)}\right)}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$=\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}+\xi_{2})(g(x)+O(\xi_{1}\varepsilon))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$\lesssim\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}+o\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}}\right)$
by the method of stationary phase applied to the phase $g$, since $g^{\prime\prime}$ is non-degenerate at the critical points ±$\pm\sqrt{\frac{1}{\beta}}$. Similarly when $\xi_{2}\gg\xi_{1}$, we have
@@
$\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$ $=\mathopen{}\mathclose{{}\left|\int_{|x|\leqslant\varepsilon}e^{-\mathrm{i}(\xi_{1}+\xi_{2})(g^{\prime}(x)+O(\xi_{2}\varepsilon))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|$
$\lesssim\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}+o\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{\xi_{1}+\xi_{2}}}}\right)$
by the method of stationary phase applied to the phase $g^{\prime}$, since $g^{\prime\prime\prime}$ is non-degenerate at the critical points 00 and ±$\pm\sqrt{\frac{3}{\beta}}$. The same argument then carries through when $\xi_{1}\sim\xi_{2}$, using the phase $g+g^{\prime}$. This yields the desired conclusion.

### v3

@@ 5 Additional proofs
$\mathopen{}\mathclose{{\left|\int_{\cup_{j}J_{j}(\theta)}e^{-\mathrm{i}\rho\phi_{\theta}(x)}\chi_{R}(x)e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}}\right|$ $\leqslant\frac{C}{\sqrt{\rho}}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left\|\chi_{R}(\cdot)e^{-\frac{(\cdot)^{2}}{2}}}}\right\|_{L^{\infty}(\mathbb{R})}+\mathopen{}\mathclose{{\left\|\mathopen{}\mathclose{{\left(\chi_{R}(\cdot)e^{-\frac{(\cdot)^{2}}{2}}}}\right)^{\prime}}}\right\|_{L^{\infty}(\mathbb{R})}}}\right)$
@@
$\leqslant\frac{C}{\sqrt{\rho}}\mathopen{}\mathclose{{\left(1+e^{-\frac{1}{2}}+\frac{\|\chi^{\prime}\|_{L^{\infty}(\mathbb{R})}}{R}}}\right),$
where $C=C(R,\beta,\delta)>0$ is independent of $\rho,\theta$. Now set $W(\theta)\coloneqq[-2R,2R]\setminus\cup_{j}J_{j}(\theta)$. By compactness and continuity, we again have $c_{W}\coloneqq\inf_{\theta\in\mathbb{S}^{1}}\inf_{x\in W(\theta)}|\phi_{\theta}^{\prime}(x)|>0$. We now look to use the method of non-stationary phase: integration by parts gives
@@
$\mathopen{}\mathclose{{\left|\int_{W(\theta)}e^{-\mathrm{i}\rho\phi_{\theta}(x)}\chi_{R}(x)e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}}\right|\leqslant\frac{1}{\rho}\int_{W(\theta)}\mathopen{}\mathclose{{\left|\frac{\mathopen{}\mathclose{{\left(\chi_{R}e^{-\frac{(\cdot^{2})}{2}}}}\right)^{\prime}(x)}{\phi_{\theta}^{\prime}(x)}-\frac{\chi_{R}(x)e^{-\frac{x^{2}}{2}}\phi_{\theta}^{\prime\prime}(x)}{(\phi_{\theta}^{\prime}(x))^{2}}}}\right|\,\mathrm{d}x$
@@
$\leqslant\frac{1}{\rho}\mathopen{}\mathclose{{\left(\frac{1}{c_{W}}\mathopen{}\mathclose{{\left\|\mathopen{}\mathclose{{\left(\chi_{R}e^{-\frac{(\cdot)^{2}}{2}}}}\right)^{\prime}}}\right\|_{L^{1}(\mathbb{R})}+\frac{M}{c_{W}^{2}}\mathopen{}\mathclose{{\left\|\chi_{R}e^{-\frac{(\cdot)^{2}}{2}}}}\right\|_{L^{1}(\mathbb{R})}}}\right),$
where $M=\sup_{x\in[-2R,2R]}\sup_{\theta\in\mathbb{S}^{1}}|\phi_{\theta}^{\prime\prime}(x)|<\infty$. All in all, both bounds yield $|\mathscr{F}\nu_{t}(\xi)|\lesssim\|\xi\|^{-\frac{1}{2}}$ for $\rho\coloneqq\|\xi\|\geqslant 1$ sufficiently large, which proves (5.7).
We now have
$\int_{\mathbb{R}^{2}}|\mathscr{F}\nu_{t}(\xi)|^{n}\,\mathrm{d}\xi\lesssim\int_{\|\xi\|\leqslant 1}1\,\mathrm{d}\xi+\int_{\|\xi\|>1}|\xi|^{-\frac{n}{2}}\,\mathrm{d}\xi,$
which is finite as long as $n>4$. By Fourier inversion, we have the desired conclusion.
@@
6 Concluding remarks
The with high probability version of the statements and the question in the higher-dimensional case remains open: we conjecture the number of modes to be $\Theta(\sqrt{\beta^{d}\log\beta})$. We also raise the question for non-Gaussian densities as well as the case of the unit sphere $\mathbb{S}^{d-1}$ with uniformly distributed samples.
Acknowledgments
The authors would like to thank Enno Mammen for useful discussion and sharing important references. We also thank Dan Mikulincer for discussions on Gaussian approximation using Edgeworth expansions, Valeria Banica for comments on the method of stationary phase, and Alexander Zimin for providing Figure˜2. We finally thank all the reviewers for their comments, which have greatly improved the quality of the paper.
Funding
B.G. was supported by a Sorbonne Emergences grant, and a gift from Google. P.R. was supported by NSF grants DMS-2022448, CCF-2106377, and a gift from Apple. Y.S. was supported by the MIT UROP and MISTI France Programs.
Borjan Geshkovski
Inria & Laboratoire Jacques-Louis Lions
Sorbonne Université
4 Place Jussieu
75005 Paris, France
e-mail: borjan.geshkovski@inria.fr
Philippe Rigollet
Department of Mathematics
Massachusetts Institute of Technology
77 Massachusetts Ave
Cambridge 02139 MA, United States
e-mail: rigollet@math.mit.edu
Yihang Sun
Department of Mathematics
Stanford University
450 Jane Stanford Way Building 380
Stanford, CA 94305, United States
e-mail: kimisun@stanford.edu